Almerinda Company estimates that total factory overhead costs will be $1,694,000 for the year. Direct labor hours are estimated to be 484,000. Required: A. For Almerinda Company, determine the predetermined factory overhead rate using direct labor hours as the activity base. B. During April, Almerinda Company accumulated 21,100 hours of direct labor costs on Job 50 and 25,200 hours on Job 51. Determine the amount of factory overhead applied to Jobs 50 and 51 in April. C. Prepare the journal entry on Apr. 30 to apply factory overhead to both jobs in April according to the predetermined overhead rate. Refer to the Chart of Accounts for exact wording of account titles.

CHART OF ACCOUNTS- Almerinda Company- General Ledger

ASSETS- 110 Cash, 121 Accounts Receivable, 125 Notes Receivable ,126 Interest Receivable, 131 Materials, 132 Work in Process ,133 Factory Overhead ,134 Finished Goods, 141 Supplies, 142 Prepaid Insurance ,143 Prepaid Expenses, 181 Land ,191 Factory, 192 Accumulated Depreciation-Factory

LIABILITIES- 210 Accounts Payable ,221 Utilities Payable ,231 Notes Payable, 236 Interest Payable ,241 Lease Payable, 251 Wages Payable ,252 Consultant Fees Payable EQUITY- 311 Common Stock, 340 Retained Earnings ,351 Dividends ,390 Income Summary

REVENUE- 410 Sales, 610 Interest Revenue

EXPENSES -510 Cost of Goods Sold, 520 Wages Expense ,531 Selling Expenses, 532 Insurance Expense ,533 Utilities Expense, 534 Office Supplies Expense ,540 Administrative Expenses, 560 Depreciation Expense-Factory ,590 Miscellaneous Expense ,710 Interest Expense

A. For Almerinda Company, determine the predetermined factory overhead rate using direct labor hours as the activity base.

? per direct labor hour B. During April, Almerinda Company accumulated 21,100 hours of direct labor costs on Job 50 and 25,200 hours on Job 51. Determine the amount of factory overhead applied to Jobs 50 and 51 in April. ?
Answer #1

a) Predetermine overhead rate = 1694000/484000 = 3.5 per labor hour

b) Overhead applied :

Job 50 = 21100*3.5 = 73850

Job 51 = 25200*3.5 = 88200

c) Journal entry

Date account and explanation debit credit
Work in process 162050
   Factory overhead 162050
(To record applied overhead)